Teal Company is constructing a building. Construction began on February 1 and was completed on December 31. Expenditures were $4,680,000 on March 1, $3,120,000 on June 1, and $7,800,000 on December 31.

Teal Company borrowed $2,600,000 on March 1 on a 5-year, 12% note to help finance construction of the building. In addition, the company had outstanding all year a 12%, 5-year, $5,200,000 note payable and an 11%, 4-year, $9,100,000 note payable. Compute avoidable interest for Teal Company. Use the weighted-average interest rate for interest capitalization purposes. (Round "Weighted-average interest rate" to 4 decimal places, e.g. 0.2152 and final answer to 0 decimal places, e.g. 5,275.)

Avoidable interest $enter the avoidable interest in dollars rounded to 0 decimal places
Answer #1
Payment Fund Used Annualised
1-march 4680000 10 months 4680000*10/12 3900000
1-june 3120000 7 months 3120000*7/12 1820000
31-Dec 7800000 0 months 7800000*0/12 0
TOTAL 5720000

2.

Loan Int
Loan specifically for building 2600000 12% 2600000*12% = 312000
Other Loans
5-year note payable 5200000 12% 5200000*12% = 624000
4-year note payable 9100000 11% 9100000*11% = 1001000
TOTAL 14300000 1625000

So, Weighted average for other loans 1625000/14300000*100 = 11.36%

Weighted Average Qualifying above (as above): 5720000

Interest on loan specifically for building (ie, 2600000): 312000

On Balance of qualifying loan @ 11.36%: (5720000-2600000)*11.36% = 354432

Total Avoidable Interest: 312000+354432 = 666432

Total Interest incurred by the company => 312000+1625000 = 1937000

Capitalize lower of Avoidable Interest OR Total Interest. So 666432.